OpenStudy (anonymous):

Solve the system of equations and choose the correct ordered pair. 4x + 5y = 22 7x + 4y = 29

OpenStudy (anonymous):

3, 2

OpenStudy (anonymous):

Multiply second equation by -5/4 -35x/4 +5y = -145/4 and add it to the first 16x-35x =88-145 -> -19x = -57 -> x = 3. Then 4y = 29-21 -> y=2. Ordered pair is (3,2)